Pure solid NaH2PO4 is dissolved in distilled water, making 100.00 ml of solution. 10.00ml of this...

  1. Pure solid NaH2PO4 is dissolved in distilled water, making 100.00 ml of solution. 10.00ml of this solution is diluted to 100.00 ml to prepare the original phosphate standard solution. Three working standard solutions are made from this by pipetting 0.8ml, 1.5ml and 3.0 ml of the original standard solution into 100.00 ml volumetric flasks. Acid and molybdate reagent are added and the solutions are diluted to 100.00 ml.

You may assume that all these absorbance measurements have already been corrected for any blank absorbance.

The absorbance of each is measured in the spectrophotometer.

Mass of NaH2PO4 (mg) 512.2
Absorbance    Standard 1 (0.8 ml)    0.1922
                          Standard 2 (1.5 ml) 0.3604
                          Standard 3 (3.0 ml) 0.7209

Calculate the following
1.
Concentration of original phosphate standard (mM)   

2. Concentration of Standard 1(mM)   

3.Concentration of Standard 2(mM)

4.Concentration of Standard 3(mM)

5.Slope of calibration line

6.A 10.00 ml volume of soft drink is treated with acid and molybdate and diluted to 100.00 ml. Its absorbance is measured. From the following data, calculate the concentration of phosphate in the soft drink (mM).

Sample absorbance       0.3826

Slope of the calibration line: 5.33      
Answer #1

1. The atomic masses are

Na: 22.990 u

H: 1.008 u

P: 30.974 u

O: 15.999 u

The gram molar mass of NaH2PO4 = (1*22.990 + 2*1.008 + 1*30.974 + 4*15.999) g/mol = 119.976 g/mol.

Mol(s) NaH2PO4 corresponding to 512.2 mg = (512.2 mg)/(119.976 g/mol)

= (512.2 mg)/(119.976 g/mol)*(1 g)/(1000 mg)

= 0.0042768 mol.

Volume of the prepared NaH2PO4 stock solution = 100.0 mL = (100.0 mL)*(1 L)/(1000 mL) = 0.100 L.

Molarity of the prepared NaH2PO4 stock solution = (0.0042768 mol)/(0.100 L)

= 0.042768 mol/L = 0.042768 M.

10.00 mL of the stock solution was diluted to 100.00 mL to prepare the original phosphate standard solution.

Use the dilution equation

M1V1 = M2V2

where M1 = molarity of the stock solution; V1 = volume of the stock solution; M2 = molarity of the original phosphate standard and V2 = final volume of the original phosphate standard.

Plug in values and obtain

(0.042768 M)*(10.00 mL) = M2*(100.00 mL)

======> M2 = (0.042768 M)*(10.00 mL)/(100.00 mL)

======> M2 = 4.2768*10-3M ≈ 4.277*10-3 M = (4.277*10-3 M)*(1000 mM)/(1 M) = 4.277 mM (ans, correct to 4 sig. figs).

2. Standard 1 is prepared by taking 0.8 mL of original phosphate standard and diluting to 100.00 mL.

Again, use the dilution equation.

M1V1 = M2V2

======> (4.277 mM)*(0.8 mL) = M2*(100.00 mL)

======> M2 = (4.277 mM)*(0.8 mL)/(100.00 mL)

======> M2 = 0.034216 mM ≈ 3.422*10-2 mM (correct to 4 sig. figs).

The concentration of standard 1 is 3.422*10-2 mM (ans).

3. Standard 2 is prepared by taking 1.5 mL of original phosphate standard and diluting to 100.00 mL.

Again, use the dilution equation.

M1V1 = M2V2

======> (4.277 mM)*(1.5 mL) = M2*(100.00 mL)

======> M2 = (4.277 mM)*(1.5 mL)/(100.00 mL)

======> M2 = 0.064155 mM ≈ 6.415*10-2 mM (correct to 4 sig. figs).

The concentration of standard 2 is 6.415*10-2 mM (ans).

4. Standard 3 is prepared by taking 3.0 mL of original phosphate standard and diluting to 100.00 mL.

Again, use the dilution equation.

M1V1 = M2V2

======> (4.277 mM)*(3.0 mL) = M2*(100.00 mL)

======> M2 = (4.277*10 mM)*(3.0 mL)/(100.00 mL)

======> M2 = 0.12831 mM ≈ 0.1281 M (correct to 4 sig. figs).

The concentration of standard 3 is 1.283*10-4 M (ans).

5. Plot absorbances vs concentrations of the standards 1, 2 and 3.

Plot of absorbance vs concentration of the standard solution.

The slope of the calibration plot is 5.6197 mM-1 (ans).